OpenStudy (anonymous):

Huck correctly solved 2(4^x) - 9^x = -6^x and arrived at the solution x = log_2/3 (h). Find the numerical value of h. Please show all work!

OpenStudy (anonymous):

take the natural log of both sides

OpenStudy (anonymous):

Natural log of both sides? I don't really see why since there is no e in the equation. Can you please explain?

OpenStudy (anonymous):

I think that since all the powers have different bases, then it doesn't matter what base you use for the logarithm at first. I anticipate change-of-base operations happening here.

OpenStudy (asnaseer):

is this the question:\[2(4^x)-9^x=-6^x\]

OpenStudy (anonymous):

Yes. THat is the problem.

OpenStudy (asnaseer):

ok, so first rephrase this in terms of prime factors to get:\[2(2^{2x})-3^{2x}=-(2*3)^x\]\[2(2^{2x})-3^{2x}=-2^x3^x\]then let:\[a=2^x\]\[b=3^x\]to get:\[2a^2-b^2=-ab\]\[2a^2-b^2+ab=0\]which can be factored to get:\[(2a-b)(a+b)=0\]can you solve from here?

OpenStudy (anonymous):

Hmm, clever. That explains the base of 2/3 for the logarithm..

OpenStudy (asnaseer):

thx CliffSedge - it took a while to realise that was the way to go :)

OpenStudy (anonymous):

2(2^x) - 3^x = 0 log_2 (x) - log_3 (x) = 0 Am I on the right path? I'm not very sure...

OpenStudy (asnaseer):

how did you get that step? you have two different log bases there?

OpenStudy (anonymous):

I plugged a and b back into the first binomial.

OpenStudy (asnaseer):

we ended up with:\[(2a-b)(a+b)=0\implies 2a=b\text{, or }a=-b\]therefore:\[2(2^x)=3^x\]or:\[2^x=-3^x\qquad\text{this has no real solutions so can be ignored}\]therefore the only real solution is:\[2(2^x)=3^x\]which can be rewritten as:\[2^{x+1}=3^x\]I'll leave you to solve for x from here...

OpenStudy (anonymous):

Alright. Thanks! Are there no solutions for the other one because there can't be a negative log?

OpenStudy (asnaseer):

correct - the other one has no real solutions.

OpenStudy (anonymous):

Just to make sure that I got it correctly: 2 * 2^x/3^x = 1 (2/3)^x = 1/2 x = log_2/3 (1/2) and since the question asks for h, the answer is 1/2?

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

Again, thanks so much! :D

OpenStudy (asnaseer):

yw :)